# Exponential quantum advantages in learning quantum observables from classical data

### v1

@@ 3 The learning problems
In the remainder of this Section we prove Theorem 1 by explicitly constructing an example of a provable classically hard time evolution problem and providing a quantum algorithm with learning guarantees. We sketch the proof ideas in the following paragraphs of the main text, while the full proofs can be found in the Appendix.
Classical hardness The hardness condition relies on the assumption $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}$ for some distribution $\mathcal{D}$. Notice that it is a widely believed assumption as, for example, the discrete logarithm problem is believed to not be in $\mathrm{HeurBPP}/\mathrm{poly}$ if the input values are sampled from the uniform distribution [BM19]. In general, by Lemma 3 in [GD23] if there exists a single $\mathcal{L}\in\text{BQP}$ that is not in HeurP/poly under some distribution, then for every BQP-complete problem there exists a distribution under which the problem is not in HeurP/poly.
Lemma 1 (Classical hardness of the time evolution learning problem).
@@
Under the assumption that there exists a distribution $\mathcal{D}^{\prime}$ such that $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$, then for any BQP-complete language there exists an associated input distribution $\mathcal{D}$ and a Hamiltonian $H_{\mathrm{hard}}$ such that no randomized polynomial-time classical algorithm $A_{c}$ is such that $h_{cl}(.)=A_{c}(T^{\bm{\alpha}},.)$ satisfies the learning condition of Def. 1 for the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ with $x$ sampled from $\mathcal{D}$.
@@
We now show an appropriate choice of the Hamiltonian $H_{\mathrm{hard}}$ such that the time evolution problem lies outside $\mathrm{HeurP}/\mathrm{poly}$ for a specific input distribution $\mathcal{D}$. Let $\mathcal{L}$ be a $\mathrm{BQP}$-complete language and $\mathcal{D}$ a distribution over input values such that $(\mathcal{L},\mathcal{D})\not\subset\mathrm{HeurP}/\mathrm{poly}$. Since $\mathcal{L}\in\mathrm{BQP}$ there exists a quantum circuit $U_{BQP}$ which decides input bitstrings $x\in\{0,1\}^{n}$ correctly on average with respect to $\mathcal{D}$. As shown more rigorously in the Appendix, measuring the $Z$ operator on the first qubit of the state $U_{BQP}\ket{\bm{x}}$ will output a positive or negative value depending if $x\in\mathcal{L}$ or not. Therefore, considering the observable $O^{\prime}=Z\otimes I\otimes...\otimes I$ the quantum model $f^{O^{\prime}}(\bm{x})=\text{Tr}[O^{\prime}\rho_{U_{BQP}}(\bm{x})]$, with $\rho_{U_{BQP}}(\bm{x})=U_{BQP}\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}_{BQP}$, correctly decides every input $\bm{x}$. We then show in the Appendix that if a classical algorithm can learn the function $g(\bm{x})=\text{Tr}[O^{\prime}\rho_{U_{BQP}}(\bm{x})]$ up to the learning condition in Def. 1, then it would be able to heuristically decide the language $\mathcal{L}$ with respect to the distribution $\mathcal{D}$. Finally, using Feynman’s idea [Fey85, Nag10] it is possible to construct a local Hamiltonian which time-evolves the initial state |$\ket{\bm{x}}$ into $U_{BQP}\ket{\bm{x}}$ in constant time [Chi04]. Considering as $H_{\mathrm{hard}}$ such constructed Hamiltonian concludes our proof. In fact, since there exists at least one concept (namely the one labeled by ${\bm{\alpha}}^{\prime}$ such that $O({\bm{\alpha}}^{\prime})=O^{\prime}$) which is not learnable by a classical algorithm, then we can conclude that for such kind of Hamiltonians the time evolution problem is not classically learnable.
@@
Quantum learnability To establish the quantum learnability of the classically hard concept class constructed above for the time evolution problem, we present a quantum algorithm directly. This algorithm accurately predicts any concept in the associated learning problem with high probability when provided with a sufficient number of training data samples. Crucially, we prove that the algorithm meets the learning condition of Definition 1 using only polynomial training samples and running in polynomial time. The central idea involves leveraging the capability of the quantum algorithm to efficiently prepare the time-evolved states $\rho_{H(\bm{x})}$, for any input local Hamiltonians $H$ and, in particular, for the hard instances of $H_{\mathrm{hard}}$ considered in our hardness result of Lemma 1.
Lemma 2 (Quantum learnability of the time evolution learning problem).
@@
There exist a quantum algorithm $A_{q}$ such that for any concept $f^{\bm{\alpha}}\in\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ considered in Lemma 1, $A_{q}$ satisfies the following. Given $n,\epsilon,\delta\geq 0$, and any training dataset $T^{{\bm{\alpha}}}_{\epsilon_{3}}$ of size
@@
$N=\mathcal{O}\left(\frac{\log(n/\delta)n}{\epsilon}\right)$ (4)
@@
with probability $1-\delta$ the quantum algorithm $A_{q}$ outputs a model $h^{*}(.)=A(T^{{\bm{\alpha}}}_{\epsilon_{3}},.)$ which satisfies the learning condition of Def. 1:
$\mathbb{E}_{\mathbf{x}\sim\mathcal{D}(\mathbf{x})}\big{[}\;|f^{{\bm{\alpha}}}(\mathbf{x})-h^{*}(\bm{x})|^{2}\;\big{]}\leq\epsilon$ (5)
@@
where $\mathcal{D}$ is the distribution the training data is sampled from.
@@
The idea for the quantum algorithm is the following. For every point $\bm{x}_{\ell}$ we construct a vector $\phi(\bm{x}_{\ell})\in[-1,1]^{m}$ of expectation values of the single Pauli strings present in $O$ on the time evolved quantum states. The model $h(\bm{x})=w\cdot\phi(\bm{x})$ is then trained on the data samples with a LASSO regression to find a $w^{*}$ so that the trained model is in agreement with the training samples, i.e. $h^{*}(\bm{x}_{\ell})=w^{*}\cdot\phi(\bm{x}_{\ell})\approx y_{\ell}$ for any $(\bm{x}_{\ell},y_{\ell})\sim\mathcal{T}^{\bm{\alpha}}_{\epsilon_{3}}$. As the $\ell_{1}$-norm of the optimal $w_{opt}={\bm{\alpha}}$ scales polynomially in $n$, imposing the constraint $||w||_{\ell_{1}}\leq B$ with $B=\mathcal{O}(n)$ in the LASSO regression will allow to obtain an error $\epsilon$ in the generalization performance using a training set of at most polynomial size. This is because the generalization error for the LASSO regression is bound linearly by $B$ [MRT18]. We write here the description of the quantum algorithm, while we leave the precise analysis of its sample and time complexity in the Appendix.
@@
    For every training point in $T^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell})\}_{i=1}^{N}$ the quantum algorithm prepares poly($n$) copies of the state $\rho_{H}(\bm{x}_{\ell})$ and computes the estimates of the expectation values $\langle P_{j}\rangle_{\ell}=\text{Tr}[\rho_{H}(\bm{x}_{\ell})P_{j}]\;\;\forall j=1,...,m$ up to a certain precision $\epsilon_{1}$. Note that $m$ scales at most polynomially in $n$ as $\{P_{j}\}_{j=1}^{m}$ are local observables.
@@
    Importantly, to meet the learning condition the optimization does not need to be solved exactly, i.e. $w^{*}={\bm{\alpha}}$. As we will make it clear in the Appendix, it is sufficient to obtain a $w^{*}$ whose training error is $\epsilon_{2}$ larger than the optimal one.
@@
3.1 Generalizations: quantum advantages for fixed inputs and for ground state problems
We showed that the concept class defined in Eq. (1) leads to a learning advantage. It is important to note that, unlike many other common learning scenario settings, the physical problem modeled by such concept class implies that the experimentalists cannot select the initial input states. However, they do possess precise control over the Hamiltonian governing the evolution of these input states. It is natural to ask if we can generalize the result and provide other cases of exponential learning advantages where the experimentalist has instead only little control over the Hamiltonian employed.
@@ 3.1 Generalizations: quantum advantages for fixed inputs and for ground state problems
Fixed inputs, unknown Hamiltonians A first example that naturally arises from the case discussed above is when the experimentalists prepare a fixed initial state and time evolves it under a Hamiltonian within a fixed family of Hamiltonians but over which they do not have control. Specifically the Hamiltonians in such a family will be labeled by some input bitstring $\bm{x}$, which for example could parameterize the strength of the coupling interactions. Each initial state will then be evolved by a different Hamiltonian in the family accordingly to the input $\bm{x}$, which comes from an unknown underlying distribution. It is easy to see that the mathematical description of such a learning problem is again defined by the concept class in Eq.(1). The only change is in the definition of $\rho_{H}(\bm{x})$, now being $\rho_{H}(\bm{x})=U(\bm{x})\ket{0}\bra{0}U(\bm{x})^{\dagger}$ with $H(\bm{x})$ the Hamiltonian associated to the quantum circuit $U(\bm{x})$ by the Feynmann construction [Fey85]. A learning advantage exists for such concept class as well, as the general definition of a language $\mathcal{L}$ in BQP implies the existence of a family of circuits $\{U(\bm{x})\}_{x}$ which correctly decides every $\bm{x}\in\mathcal{L}$.
@@
Ground state problem As our next example we consider the task of predicting ground state properties of local Hamiltonians. Here, the states to be measured are no longer time evolved quantum states but rather ground states of local Hamiltonians. Specifically, the experimentalist is capable of preparing ground states of input $k$ local Hamiltonians, which once again belong to a family of Hamiltonians. However, they lack control over the coupling parameters that, instead, are random values drawn from an underlying distribution. This situation occurs, for example, in the context of the random Ising [Nat98] or random Heisenberg model [OS01]. The mathematical formalization of such a learning problem is the following. Consider a family of parametrized local Hamiltonians =$\mathcal{H}=\{H(\bm{x})\;\;|\;\;\bm{x}\in\{0,1\}^{n}\}$. Let us define the concept class for the ground state learning problem similarly to the time evolution case of Eq. (1), where now the unknown observable $O({\bm{\alpha}})$ is measured on the states $\rho_{H}(\bm{x})$, which correspond to the ground states of the Hamiltonians $H(\bm{x})\in\mathcal{H}$, then we define:
$\mathcal{F}^{\mathcal{H},O}_{\mathrm{g.s.}}$ =$=\{f^{\bm{\alpha}}(\bm{x})\in\mathbb{R}\;\;|\;\mathbf{{\bm{\alpha}}}\in[-1,1]^{m}\}$ (7)
with $\text{with }f^{\bm{\alpha}}(\bm{x}):\;\;$ $\;\bm{x}\in\{0,1\}^{n}\rightarrow f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]\ \;\;\&\;\;O({\bm{\alpha}})=\sum_{i=1}^{m}{\bm{\alpha}}_{i}P_{i}.$ (8)
@@
Considering the training data $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell}\approx_{\epsilon_{3}}f^{\bm{\alpha}}(\bm{x}_{\ell}))\}_{\ell=1}^{N}$, the learning condition remains the same as in Definition 1. From the hardness result of Lemma 1, we obtain the following Theorem
@@
Under the assumption that there exists a distribution $\mathcal{D}^{\prime}$ such that $\mathrm{BQP}\not\subset\mathrm{HeurP}^{\mathcal{D}^{\prime}}/\mathrm{poly}$, then for any BQP-complete language there exists an associated distribution $\mathcal{D}$ over the input values $x\in\{0,1\}^{n}$ and a family of Hamiltonians $\mathcal{H}_{\mathrm{hard}}$ such that no classical algorithm can learn the ground state problem, formalized by learning the concept class $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ in the sense of Def.1 given data $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{3}}=\{(\bm{x}_{\ell},y_{\ell}\approx f^{\bm{\alpha}}(\bm{x}))\}_{\ell=1}^{N}$ sampled from $\mathcal{D}$. However, there exists a quantum algorithm which learns $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ under any distribution $\mathcal{D}$.
@@
The existence of a class of Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ for which the ground state problem is classically hard to learn is guaranteed by the argument above regarding the hardness of time evolution case. The structure of the proof is exactly the same of the proof for Theorem 1 rigorously written in the Appendix. The only missing step for the ground state version is that now the states $\rho_{H}(\bm{x})$ we are considering in Eq. (7) are ground state of Hamiltonians $H(\bm{x})$ and not time evolved states. However, using the Kitaev’s construction [KSV02, KR03], it is possible to create for any $\bm{x}\in\{0,1\}^{n}$ and $U$ a local Hamiltonian $H(\bm{x})$ such that its ground state will have a large overlap with $U\ket{\bm{x}}$, with $U$ an arbitrary quantum circuit with a polynomial depth. This completes our proof of classical hardness as we consider the family of Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ to exactly be the set of such $\{H(\bm{x})\}_{x}$ with $U$ implementing a BQP-complete computation [^2] More rigorously it means that for every $n$, we consider $U=U^{n}_{BQP}$ from the family $\{U^{n}_{BQP}\}_{n}$ of quantum circuits which correctly decide a BQP-complete language $\mathcal{L}$. Same as before, it is still the case that there is at least one concept which can not be evaluated by polynomially sized classical circuits. Finally, also the quantum algorithm with learning guarantees for the concept class $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ closely follows Algorithm 1. The only missing point to prove here is that the ground states $\rho_{\mathcal{H}_{\mathrm{hard}}}(\bm{x})$ are easily preparable on a quantum computer. Recall that the class of hard Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ considered in the hardness result of the ground state problem are the one derived from the Kitaev’s circuit-to-Hamiltonian construction from a $\mathrm{BQP}$-complete circuit. It is well known that those Hamiltonians present a $\Omega(1/\text{poly}(n))$ gap (in contrast to the classically learnable Hamiltonians in [HKT+22b]) and it is possible to construct the ground state $\rho_{\mathcal{H}_{\mathrm{hard}}}$ from the description of the corresponding Kitaev’s Hamiltonian, known to the learner through the description of the concept class and the input $\bm{x}$. This then concludes our proof for the quantum learnability of $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$, thus completing the entire proof of Theorem 2
Notice that if our interest is only about the classical hardness or the quantum learnability of the ground state problem we can consider concept classes associated to more general Hamiltonian families than the Kitaev’s family $\mathcal{H}_{\mathrm{hard}}$ used in our provable advantage result. Specifically, classical hardness for the time evolution problem is guaranteed for a number of physically motivated Hamiltonians, such as the Heisenberg model whose dynamics can efficiently be simulated on a quantum computer. Notice that for the ground state version of the learning problem considering generic local Hamiltonian might make the problem quantum not solvable. It is in fact known that finding ground states of even two local Hamiltonians, as for example the Heisenberg model, is QMA-hard. Concept classes for which the quantum learnability condition is met are the ones associated to any family of Hamiltonian $\mathcal{H}_{\mathrm{quantum}}$ for which a quantum algorithm can prepare the corresponding ground state given the Hamiltonian description. In that regard a learning speed-up can still be achieved if we consider providing to the algorithm guiding states for any family of local Hamiltonians, this in fact will make the problem $\mathrm{BQP}$-complete [GHGM22].
@@
We note that the ground state learning problem defined by the class in Eq.(7) very much resembles the ML problem studied in the works [HKT+22b] and [LHT+24] where the authors showed that a classical algorithm could solve the task. As their result holds for any local Hamiltonian with constant gap, this gives a constraint on the family of local Hamiltonians $\mathcal{H}$ one would like to consider in order to prove a quantum advantage in learning.

### v2

@@ IV The learning problems
In the remainder of this Section we prove Theorem 1 by explicitly constructing an example of a provable classically hard time-evolution problem and providing a quantum algorithm with learning guarantees. We sketch the proof ideas in the following paragraphs of the main text, while the full proofs can be found in the Appendix A and B.
IV.1 Classical hardness
Our classical hardness results are based on the assumption $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$. As previously stated, this is different from previous works [11], where classical intractability was established by considering the distributional version of complexity classes. Specifically, the results in [11] explicitly assume the existence of an input distribution $\mathcal{D}_{\mathrm{hard}}$ for which no classical algorithm can even approximate, in the sense of Eq. 1, a target function that computes $\operatorname{\mathsf{BQP}}$-complete languages. In contrast, this work considers the weaker assumption that $\operatorname{\mathsf{BQP}}$-hard languages cannot be correctly decided on every input by polynomial-sized classical circuits. While the assumption considered in [11] implies the assumption considered in this paper, the reverse is not generally true. To achieve classical hardness through the more natural and studied assumption $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$, we require a stronger but arguably more natural learning condition in Def. 4 than what present in [11]. Namely, we work in the distribution free PAC framework and require the learning algorithm to succeed for any input distribution. In this way, we can use well-known results from classical learning theory that link the hardness of learning to the hardness of computing a target function. The full proof of classical hardness, as well as quantum learnability, for the learning task can be found in Appendix A and Appendix B respectively. To prove classical hardness is enough to show that there exists a local Hamiltonian $H_{\mathrm{hard}}$ for which the concepts considered in our learning problem can decide $\operatorname{\mathsf{BQP}}$ languages. In fact, by the results in [23] from classical learning theory, if a target concept is learnable in the sense of Def. 1, then there exists a polynomial size classical circuit which efficiently evaluates it on every input point. In particular, this implies that the concept can be computed in $\mathsf{P/poly}$. Therefore, if the concept class $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$ is classically learnable, then $\operatorname{\mathsf{BQP}}\subseteq\mathsf{P/poly}$. As a final note, we observe that although our learning problem is defined over real-valued functions, in our hardness result, the specific Hamiltonian $H_{\mathrm{hard}}$ and the associated observables ensure that the target concept has a binary output.
Theorem 2 (Classical hardness of the time-evolution learning problem).
@@ IV.1 Classical hardness
For any $\operatorname{\mathsf{BQP}}$-complete language, there exists a Hamiltonian $H_{\mathrm{hard}}$ such that no randomized polynomial-time classical algorithm $A_{c}$ satisfies the learning condition of Def. 4 for the concept class $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$, unless $\operatorname{\mathsf{BQP}}\subseteq\mathsf{P/poly}$.
Proof sketch.
@@
The detailed proof of Theorem 2 can be found in the Appendix A, here we give the overall proof strategy. Let $\mathcal{L}$ be an arbitrary $\operatorname{\mathsf{BQP}}$ language. Since $\mathcal{L}\in\operatorname{\mathsf{BQP}}$ there exists a quantum circuit $U_{BQP}$ which decides input bitstrings $\bm{x}\in\{0,1\}^{n}$ correctly on average with respect to $\mathcal{D}$. As shown more rigorously in the Appendix A, measuring the $Z$ operator on the first qubit of the state $U_{BQP}\ket{\bm{x}}$ will output a positive or negative value depending if $\bm{x}\in\mathcal{L}$ or not. Therefore, for the observable $O^{\prime}=Z\otimes I\otimes...\otimes I$, the quantum model $f^{O^{\prime}}(\bm{x})=\text{Tr}[O^{\prime}\rho_{U_{BQP}}(\bm{x})]$, with $\rho_{U_{BQP}}(\bm{x})=U_{BQP}\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}_{BQP}$, correctly decides every input $\bm{x}$. Finally, using Feynman’s idea [24, 25] it is possible to construct a local Hamiltonian which time-evolves the initial state |$\ket{\bm{x}}$ into $U_{BQP}\ket{\bm{x}}$ in constant time [26]. Denote this constructed Hamiltonian as $H_{\mathrm{hard}}$. Then the concept $\bm{\alpha^{\prime}}$, associated to the observable $O({\bm{\alpha}}^{\prime})=O^{\prime}$, implements a $\operatorname{\mathsf{BQP}}$ computation. The final step is a result in [23] which guarantees that if a function is learnable under any distribution in the sense of Def. 1, then there exists a polynomial size circuit which correctly evaluates it on every input $\bm{x}$. We provide a more detailed explanation of the crucial result contained in [23] and its consequences in Appendix A. This concludes the proof, as if $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ was learnable, then there would be a polynomial size circuit which evaluates the concept $f^{\alpha^{\prime}}$. Thus any $\operatorname{\mathsf{BQP}}$ language could be decided in $\mathsf{P/poly}$. ∎
IV.2 Quantum learnability
@@
To establish the quantum learnability of the classically hard concept class constructed above for the time-evolution problem, we present a quantum algorithm directly. This algorithm accurately predicts any concept in the associated learning problem with high probability when provided with a sufficient number of training data samples. Crucially, we prove that the algorithm meets the learning condition of Def. 4 using only polynomial training samples and running in polynomial time. The central idea involves leveraging the capability of the quantum algorithm to efficiently prepare the time-evolved states $\rho_{H(\bm{x})}$, for an input local Hamiltonians $H$ and, in particular, for the hard instances of $H_{\mathrm{hard}}$ considered in our hardness result of Lemma 2.
Theorem 3 (Quantum learnability of the time-evolution learning problem).
@@
There exist an efficient quantum algorithm $A_{q}$ such that for any concept $f^{\bm{\alpha}}\in\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ considered in Lemma 2, $A_{q}$ satisfies the following. Given $n,\epsilon,\delta\geq 0$, and any training dataset $T^{{\bm{\alpha}}}_{\epsilon_{2}}$, with $\epsilon_{2}\leq\epsilon$, of size
@@
$N=\mathcal{O}\left(\frac{\log(\mathrm{poly}(n)/\delta)\mathrm{poly}(n)}{\epsilon^{2}}\right)$ (5)
@@
with probability $1-\delta$ the quantum algorithm $A_{q}$ outputs a model $h^{*}(.)=A(T^{{\bm{\alpha}}}_{\epsilon_{2}},\epsilon,\delta,.)$ which satisfies the learning condition of Def. 4:
$\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\big{[}\;|f^{{\bm{\alpha}}}(\mathbf{x})-h^{*}(\bm{x})|^{2}\;\big{]}\leq\epsilon$ (6)
@@
where $\mathcal{D}$ is any arbitrary distribution from which the training data is sampled.
@@
The idea for the quantum algorithm is the following. For every point $\bm{x}_{\ell}$ we construct a vector $\phi(\bm{x}_{\ell})\in[-1,1]^{m}$ of expectation values of the single Pauli strings present in $O$ on the time evolved quantum states. The model $h(\bm{x})=w\cdot\phi(\bm{x})$ is then trained on the data samples with a LASSO regression to find a $w^{*}$ so that the trained model is in agreement with the training samples, i.e. $h^{*}(\bm{x}_{\ell})=w^{*}\cdot\phi(\bm{x}_{\ell})\approx y_{\ell}$ for any $(\bm{x}_{\ell},y_{\ell})\sim\mathcal{T}^{\bm{\alpha}}_{\epsilon_{2}}$. As the $\ell_{1}$-norm of the optimal $w_{opt}={\bm{\alpha}}$ scales polynomially in $n$, imposing the constraint $||w||_{\ell_{1}}\leq B$ with $B=\mathcal{O}(\mathrm{poly}(n))$ in the LASSO regression will allow to obtain an error $\epsilon$ in the generalization performance using a training set of at most polynomial size. This is because the generalization error for the LASSO regression is bounded linearly by $B$ [17]. The description of the quantum algorithm can be found in Algorithm 1, while we leave the precise analysis of its sample and time complexity in the Appendix B.
@@
    For every training point in $T^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{2}}=\{(\bm{x}_{\ell},y_{\ell})\}_{i=1}^{N}$ the quantum algorithm prepares poly($n$) copies of the state $\rho_{H}(\bm{x}_{\ell})$ and computes the estimates of the expectation values $\langle P_{j}\rangle_{\ell}=\text{Tr}[\rho_{H}(\bm{x}_{\ell})P_{j}]\;\;\forall j=1,...,m$ up to a certain precision $\epsilon_{1}$. Note that $m$ scales at most polynomially in $n$ as $\{P_{j}\}_{j=1}^{m}$ are local observables.
@@
    Importantly, to meet the learning condition the optimization does not need to be solved exactly, i.e. $w^{*}={\bm{\alpha}}$. As we will make it clear in the Appendix B, it is sufficient to obtain a $w^{*}$ whose training error is $\epsilon_{2}$ larger than the optimal one.
@@
IV.3 Generalizations: quantum advantages for fixed inputs and for ground state problems
We showed that the concept class defined in Eq. (3) leads to a learning advantage. In this case, the physical problem modeled by $\mathcal{F}_{\mathrm{evolved}}^{H,O}$ assumes the initial input states are drawn from an underlying distribution, while there is precise control over the Hamiltonian governing their evolution. This raises the question of whether learning advantages can be extended to other scenarios, modeled by cases with limited control over the Hamiltonian. We argue that the following scenarios can represent practical examples in Section VI.
@@ IV.3 Generalizations: quantum advantages for fixed inputs and for ground state problems
Fixed inputs, unknown Hamiltonians A first example that naturally arises from the case discussed above is when a fixed initial state is prepared and time-evolved under a Hamiltonian from a fixed family of Hamiltonians, over which there is no full control. Specifically the Hamiltonians in such a family will be labeled by some input bitstring $\bm{x}$, which for example could parameterize the strength of the coupling interactions. Each initial state will then be evolved by a different Hamiltonian in the family accordingly to the input $\bm{x}$, which comes from an unknown underlying distribution. It is easy to see that the mathematical description of such a learning problem is again defined by the concept class in Eq.(3). The only change is in the definition of $\rho_{H}(\bm{x})$, now being $\rho_{H}(\bm{x})=U(\bm{x})\ket{0}\bra{0}U(\bm{x})^{\dagger}$ with $H(\bm{x})$ the Hamiltonian associated to the quantum circuit $U(\bm{x})$ by the Feynman construction [24]. A learning advantage exists for such concept class as well, as the general definition of a language $\mathcal{L}$ in BQP implies the existence of a family of circuits $\{U(\bm{x})\}_{x}$ which correctly decides every $\bm{x}\in\mathcal{L}$.
@@
Ground state problem As the next example, we consider predicting ground state properties of local Hamiltonians. Here, the states to be measured are the ground states of local Hamiltonians, rather than time-evolved quantum states. Specifically, ground states of input $k$-local Hamiltonians, which belong to a family of Hamiltonians, are prepared (Section VI discusses when this is feasible on a quantum computer). However, there is no full control over the coupling parameters, which are instead random values drawn from an underlying distribution. This situation is close, for example, to the case of the random Ising [27] or random Heisenberg model [28]. The mathematical formalization of such a learning problem is the following. Consider a family of parametrized local Hamiltonians =$\mathcal{H}=\{H(\bm{x})\;\;|\;\;\bm{x}\in\{0,1\}^{n}\}$. Let us define the concept class for the ground state learning problem similarly to the time-evolution case of Eq. (3), where now the unknown observable $O({\bm{\alpha}})$ is measured on the states $\rho_{H}(\bm{x})$, which correspond to the ground states of the Hamiltonians $H(\bm{x})\in\mathcal{H}$, then we define:
$\mathcal{F}^{\mathcal{H},O}_{\mathrm{g.s.}}=\{f^{\bm{\alpha}}(\bm{x})\in\mathbb{R}\;\;|\;\mathbf{{\bm{\alpha}}}\in[-1,1]^{m}\}$ (8)
with: $f^{\bm{\alpha}}(\bm{x}):\;\;\;\bm{x}\in\{0,1\}^{n}\rightarrow f^{\bm{\alpha}}(\bm{x})=\text{Tr}[\rho_{H}(\bm{x})O(\mathbf{{\bm{\alpha}}})]$
$O({\bm{\alpha}})=\sum_{i=1}^{m}{\bm{\alpha}}_{i}P_{i}.$
@@
Considering the training data $\mathcal{T}^{\mathbf{{\bm{\alpha}}}}_{\epsilon_{2}}=\{(\bm{x}_{\ell},y_{\ell}\approx_{\epsilon_{2}}f^{\bm{\alpha}}(\bm{x}_{\ell}))\}_{\ell=1}^{N}$, the learning condition remains the same as in Definition 4. From the hardness result of Lemma 2, we obtain the following Theorem
@@
For any $\operatorname{\mathsf{BQP}}$-complete language there exists a family of Hamiltonians $\mathcal{H}_{\mathrm{hard}}$ such that no classical algorithm can learn the ground state problem, formalized by learning the concept class $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ in the sense of Def.4, unless $\operatorname{\mathsf{BQP}}\subseteq\mathsf{P/poly}$. However, there exists a quantum algorithm which learns $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ under any input distribution $\mathcal{D}$.
@@
The existence of a class of Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ for which the ground state problem is classically hard to learn is guaranteed by the argument above regarding the hardness of time-evolution case. The structure of the proof is exactly the same of the proof for Theorem 1 rigorously written in the Appendix A. The only missing step for the ground state version is that now the states $\rho_{H}(\bm{x})$ we are considering in Eq. (8) are ground state of Hamiltonians $H(\bm{x})$ and not time evolved states. However, using the Kitaev’s construction [29, 30], it is possible to create for any $\bm{x}\in\{0,1\}^{n}$ and $U$ a local Hamiltonian $H(\bm{x})$ such that its ground state will have a large overlap with $U\ket{\bm{x}}$, with $U$ an arbitrary quantum circuit with a polynomial depth. This completes our proof of classical hardness as we consider the family of Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ to exactly be the set of such $\{H(\bm{x})\}_{x}$ with $U$ implementing a $\operatorname{\mathsf{BQP}}$-complete computation  [^2] More rigorously it means that for every $n$, we consider $U=U^{n}_{BQP}$ from the family $\{U^{n}_{BQP}\}_{n}$ of quantum circuits which correctly decide a $\operatorname{\mathsf{BQP}}$-complete language $\mathcal{L}$. Same as before, it is still the case that there is at least one concept which can not be evaluated by polynomially sized classical circuits. Finally, also the quantum algorithm with learning guarantees for the concept class $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$ closely follows Algorithm 1. The only missing point to prove here is that the ground states $\rho_{\mathcal{H}_{\mathrm{hard}}}(\bm{x})$ are easily preparable on a quantum computer. Recall that the class of hard Hamiltonian $\mathcal{H}_{\mathrm{hard}}$ considered in the hardness result of the ground state problem are the one derived from the Kitaev’s circuit-to-Hamiltonian construction from a $\operatorname{\mathsf{BQP}}$-complete circuit. It is well known that those Hamiltonians present a $\Omega(1/\text{poly}(n))$ gap (in contrast to the classically learnable Hamiltonians in [9]) and it is possible to construct the ground state $\rho_{\mathcal{H}_{\mathrm{hard}}}$ from the description of the corresponding Kitaev’s Hamiltonian, known to the learner through the description of the concept class and the input $\bm{x}$. This then concludes our proof for the quantum learnability of $\mathcal{F}^{\mathcal{H}_{\mathrm{hard}},O}_{\mathrm{g.s.}}$, thus completing the entire proof of Theorem 4
It is important to note that although our separation results, both for the time-evolution and the ground state versions of the problem, are proven using Kitaev’s Hamiltonians, analogous results hold for a broader and more physical class of Hamiltonians. Specifically, since we rely on the assumption that $\operatorname{\mathsf{BQP}}\not\subseteq\mathsf{P/poly}$, any quantum process that cannot be simulated by polynomial-sized circuits gives rise to a learning problem of the kind we introduced, with a provable quantum learning advantage. In Section VI we list examples of physical Hamiltonians to which our results apply, covering both the time-evolution and ground state versions of the problem.
@@
We note that the ground state learning problem defined by the class in Eq. (8) closely resembles the machine learning problem studied in [9] and [10], where the authors demonstrated that a classical algorithm could solve the task, provided the Hamiltonian has a constant gap. Since the Hamiltonians considered in our results have a polynomially decaying gap, this imposes a constraint on the family of local Hamiltonians $\mathcal{H}$ that one would need to consider in order to prove a quantum advantage in learning.